Use below information for Questions 7 to 10:

Jane Botosan operates a bed and breakfast hotel in a resort area near Lake Michigan. Depreciation on the hotel is $60,000 per year. Jane employs a maintenance person at an annual salary of $41,000 and a cleaning person at an annual salary of $24,000. Real estate taxes are $10,600 per year. The rooms rent at an average price of $127 per person per night including breakfast. Other costs are laundry and cleaning service at a cost of $10.6 per person per night and the cost of food which is $5.3 per person per night.

Q-7) Determine the sales revenue Jane needs to break even.

Q-8) Determine the number of rentals Jane needs to earn a target net income of $50,000.

Q-9) Jane is considering changing the business strategy. She considers installing additional cleaning machines which will decrease laundry and cleaning service per person per night by $3. However, fixed depreciation cost will increase by $4,030. Determine the number of rentals Jane needs to break even if the changes are made.

Q-10) Determine the number of rentals at which Jane would be indifferent between the current and proposed business models. (Hint: Consider net income figures.)
Answer #1
Answer to Question 7 :
Fixed costs = $60,000+$41,000+$24,000+$10,600 = $135,600
Contribution per unit = $127-$10.6-$5.3 = $111.1 p.u
Contribution margin ratio = 111.1/127 = 87.48%
Sales revenue to break even = Fixed costs/Contribution margin ratio = 135,600/87.48% = $1,55,006
Answer to Question 8 :
Number of rentals to be sold to earn net income of $50,000 = (135,600+50,000)/111.1 = 1,670 units(rounded off)
Answer to Question 9 :
New fixed costs = 135,600+4,030 = $139,630
New Contribution per unit = 127-7.6-5.3 = $114.1 p.u
Rental units to break-even = 139,630/114.1 = 1,224 units (rounded off)
Answer to Question 10 :
Net Income = Total contribution - Fixed costs
Let number of rentals at which Jane would be indifferent be "x".
Now, when x number of rentals are sold, Jane would get equal net income under both cases
Hence, we can calculate x by using the following equation :
111.1x - 135,600 = 114.1x - 139,630
3x = 139,630-135,600
3x = 4030
x = 1,343 units (rounded off)
Therefore, number of rentals at which Jane would be indifferent among both options = 1,343 units